A radioactive element has an initial count rate of 1400 counts per second as measured by proportional counter and doors falls to 1800 counts per minute in 15 hours

Calculate the half-life of the radioactive elements

if the initial number of atoms in their sample of this element is 4.5 * 10 raise to power 9 how many albums has decayed in 24 hours


1
Expert's answer
2022-03-24T00:06:04-0400

Half life= {1800-1400}/15× 60× 60= 7.4 secs

"N\/ N\u2070= exp(-\u00b6t)"

¶= 0.623/7.4= 0.084

N= 4.5× 10^9exp(-0.084×24)

N= 2.4× 10^8